OpenStudy (anonymous):

Just solve. 1/2 + 3/t = 5/8

OpenStudy (cwrw238):

aren'ytyou missing the variable?

OpenStudy (cwrw238):

* aren't you

OpenStudy (anonymous):

I don't know why I had 4 sorry, it's t.

OpenStudy (zzr0ck3r):

t is the variable.

OpenStudy (zzr0ck3r):

multiply every term by t and what do you have?

OpenStudy (cwrw238):

multiply each term by the gcf of 2 , t and 8

OpenStudy (anonymous):

So far I have .\[2x(\frac{ 1 }{ 2 } + \frac{ 3 }{ t }) = 2x \frac{ 5 }{ 8 }\]

OpenStudy (zzr0ck3r):

no x's

OpenStudy (zzr0ck3r):

x=t

OpenStudy (anonymous):

oh alright. then that equation except with the t's.

OpenStudy (zzr0ck3r):

its just a "place holder"

OpenStudy (zzr0ck3r):

good, now distribute

OpenStudy (anonymous):

I then got 2x(1/2) + 2x(3/4) = 2x(5/8) and in the first parenthesis crossed out both 2's and in the second set of parenthesis crossed out the variables. Now what do I do with the third set?

OpenStudy (zzr0ck3r):

that's 10x/8

OpenStudy (anonymous):

Alright. Thanks.

OpenStudy (anonymous):

Help with another?

OpenStudy (anonymous):

Wait, after 10x8 I got 1t + 6t = 10x8 is that the answer?

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

1/2 + 3/x = 5/8 x(1/2+3/x)=5x/8 x/2+3=5x/8 multiply everything by 8 to kill the 8 and 2 8(x/2+3)=8*5x/8 4x+24=5x x=24